# Altermagnetism-induced non-collinear superconducting diode effect and unidirectional superconducting transport

### v1

@@ Abstract
Given their time-reversal-breaking yet inversion-symmetric character, and the absence of the orbital effects due to a net magnetization, altermagnets provide a natural platform for exploring their interplay with superconductivity [46, 47, 48]. The intrinsic momentum-dependent spin polarization in altermagnets induces a mismatch between spin-up and spin-down Fermi surfaces, favoring the finite-momentum Cooper pairing even in the absence of external fields [49, 50]. Recent theoretical works [51, 52, 53, 54, 55] have investigated the stability of such pairing near the SC transition, typically by solving the linearized gap equation. A fully self-consistent microscopic analysis was performed by Hong et al. [27], who systematically compared conventional BCS pairing, chiral $p$-wave states, and FF pairing under comparable altermagnetic conditions. Their results revealed that, within a narrow but realistic range of altermagnetic strength, the FF state can become energetically favorable without requiring external magnetic fields or fine-tuned SOC.
Motivated by these developments [50, 49, 27], here we investigate directional SC transport in conventional $s$-wave SC thin films proximity-coupled to a $d$-wave altermagnet, using a fully microscopic and self-consistent framework. As the altermagnetic strength increases from zero, we find that the critical current in the BCS regime evolves from being isotropic to strongly anisotropic with a fourfold ($C_{4}$) symmetry, reflecting the underlying crystalline symmetry of the altermagnetic spin texture. Thus, the directions of maximal and minimal critical currents are not antiparallel, indicating that the non-reciprocity deviates from the simple single-axis diode behavior and instead reflects a non-collinear SC diode effect.
@@
[…]
@@
Diagonalizing this Hamiltonian, one obtains the gap equation:
@@
where the quasiparticle energy spectrum is given by $E_{\bf k}^{\pm}=\hbar{\bf v}_{\bf k}\cdot{\bf q}+h_{\bf k}\pm{E_{\bf k}}$ with ${\bf v}_{\bf k}=\hbar{\bf k}/m$ and $E_{\bf k}=\sqrt{\Xi_{k}^{2}+\Delta_{\bf q}^{2}}$ [27, 50]. Here, $\Xi_{\bf k}=\xi_{\bf k}+2\gamma_{d}(k_{x}q_{x}-k_{y}q_{y})$ and $f(x)$ is the Fermi-Dirac distribution function. The quasiparticle energies here are tilted by the finite CM momentum and exhibit particle–hole asymmetry. As a result, at zero temperature, the integrand in the gap equation becomes nonzero only when $E_{\bf k}^{+}>0$ and $E_{\bf k}^{-}<0$. Thus, SC anomalous pairing correlations vanish in momentum-space regions where either $E_{\bf k}^{+}<0$ or $E_{\bf k}^{-}>0$, as $f(E_{\mathbf{k}}^{+})-f(E_{\mathbf{k}}^{-})\equiv 0$. These regions are referred to as unpairing regions [16, 20, 64], where electrons no longer contribute to Cooper pairing and instead behave as normal-state excitations.
The SC free energy of the system is given by
$\mathcal{F}^{s}_{\mathbf{q}}=-\frac{1}{\beta}\sum_{\mathbf{k},\lambda=\pm}\ln\big{(}1+e^{-\beta\lambda E_{\mathbf{k}}^{\lambda}}\big{)}-\sum_{\mathbf{k}}\left(E_{\bf k}-\Xi_{\mathbf{k}}\right)-\frac{|\Delta_{\mathbf{q}}|^{2}}{U},$ (4)
@@
where $\beta=1/(k_{B}T)$. The equilibrium SC state is determined by self-consistently solving the gap equation and minimizing the free energy $\mathcal{F}^{s}_{\bf q}$ with respect to the pairing momentum ${\bf q}$. This procedure allows us to determine the optimal CM momentum ${\bf q}_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as a function of system parameters such as the altermagnetic strength and temperature. To investigate the transport properties, we introduce an additional SC momentum ${\bf p}_{s}$ [65, 66, 67, 68, 64, 69, 60, 17, 70, 71], corresponding to a uniform phase gradient, superflow, or external vector potential. The resulting supercurrent is given by ${\bf j}_{s}=(e/m)n_{s}{\bf p}_{s}$ [65, 66, 67, 68, 64, 69, 60, 72], where $n_{s}$ is the superfluid density [73], which approaches the total electron density $n$ in the low-temperature limit [69, 63, 20, 65, 17, 74, 75]. To determine the critical current, we gradually increase the magnitude of ${\bf p}_{s}$ along a fixed direction and solve the gap equation self-consistently under the total momentum ${\bf q}_{0}+{\bf p}_{s}$. The SC gap collapses at a critical value ${\bf p}_{s,c}$, signaling the pair-breaking threshold and marking the onset of dissipation.
@@
For comparison, we also consider a conventional FFLO model with an isotropic Zeeman spin splitting arising from ferromagnetic proximity, modeled as $h_{\bf k}=\gamma_{s}k^{2}$. This serves to isolate the role of the anisotropic, crystal-symmetry-protected spin splitting characteristic of altermagnets, and to demonstrate how it gives rise to robust finite-momentum pairing and directional SC diode effects.
@@
Figure 2: Simulated order-parameter (OP) results for the optimal CM momentum $q_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as functions of the spin-splitting strength for (a) ferromagnetic and (b) $d$-wave altermagnetic proximity. Insets: the $\Delta_{\bf q}$ landscape plotted in $q_{x}$$q_{y}$ plane, which reflects the inverse of the free-energy profile. The plotted $q$ range extends up to $0.8q_{0}$ in each direction. $T=0.1~{}$K. Normalized parameters used here are: the zero-splitting gap $\Delta_{0}=2~{}$meV, coherence length $\xi_{0}=v_{F}/\Delta_{0}$, and splitting unit $\gamma_{0}=\Delta_{0}/k_{F}^{2}$.
@@
Phase diagram.— As a specific example, we consider a typical SC thin film, NbN, which has a SC gap $\Delta_{0}\sim 2~{}\text{meV}$ at $T\sim 0$ and a critical SC temperature $T_{c}\sim 12$-15 K [63, 76, 77]. Due to the Cooper instability of interacting fermions underlying the BCS theory [17, 61, 60, 78], the momentum summation for the pairing electrons is restricted such that the spin-up electron with momentum ${\bf k+q}$ and the spin-down electron with momentum ${\bf-k+q}$ both lie within their respective Debye momentum shells around the Fermi surface, enforced by the spin-dependent condition $|\xi_{s{\bf k+q}}+sh_{s{\bf k+q}}|\leq\omega_{D}$. It is noted that the CM momentum dependence of $\Delta_{\bf q}$ directly reflects the inverse of the free-energy landscape since the condensation energy satisfies $\mathcal{F}^{s}_{\mathbf{q}}-\mathcal{F}^{n}_{\mathbf{q}}\propto-\Delta_{\bf q}^{2}$, offering an efficient means for following analysis of the free-energy minimum.
@@
[…]
@@
For an altermagnetic exchange field, the evolution of the critical current exhibits qualitatively distinct behavior. As the splitting increases from zero, the BCS state [Fig. 3(c)] gradually loses its isotropy, and the angular profile of $J_{c}(\delta\phi)$ starts to develop the pronounced oscillations. With increasing $\gamma_{d}$, the critical current along the spin-splitting axes (i.e., the $x$ and $y$ directions: $\delta\phi=\pm\pi/4$ and ±$\pm 3\pi/4$) is progressively suppressed, while the critical current along the nodal directions of the splitting (i.e., the diagonal directions: $\delta\phi=0,\pi$ and ±$\pm\pi/2$) is enhanced. As a result, $J_{c}(\delta\phi)$ acquires a pronounced fourfold ($C_{4}$) anisotropy, directly reflecting the $d$-wave character of the altermagnetic spin texture and corresponding to the $C_{4}$-symmetric, square-like shape of the $\Delta_{\bf q}$ landscape below the critical value [see the insets of Fig. 2(b)]. Thus, the maxima of $J_{c}(\delta\phi)$ occur along the diagonal directions, while the minima lie along $x$ and $y$ axes. This nontrivial angular misalignment between the directions of maximal and minimal critical currents gives rise to a non-collinear SC diode effect, where the non-reciprocity of the supercurrent is not aligned with a single transport direction, but instead emerges from distinct crystal axes. This non-collinear SC diode effect is robust, as it originates from the crystal-symmetry-protected nature of the altermagnetic spin splitting, enabling intrinsically stable and symmetry-guided rectification of supercurrents.
@@
With further increase of $\gamma_{d}$, following the onset of the FF phase, a finite CM momentum ${\bf q}_{0}$ develops along one of the nodal directions of the altermagnetic spin splitting, and directly breaks SC transport along the opposite direction. This asymmetry occurs because ${\bf q}_{0}$ resides near the inner boundary of the gap landscape $\Delta_{\bf q}$, making the system particularly sensitive to supercurrents applied in the $-{\bf q}_{0}$ direction, which easily push the total momentum beyond the de-pairing threshold and trigger SC breakdown. As a result, only three residual peaks remain in the critical-current profile, aligned with the other nodal directions ($\delta\phi=0,\pm\pi/2$), since in this early FF regime, the $\Delta_{\bf q}$ landscape still spans a continuous connected region in ${\bf q}$ space [see the inset of Fig. 2(b)], although its global maximum has already shifted away from the origin to ${\bf q}_{0}$. Notably, as $\gamma_{d}$ increases further, the critical-current peaks perpendicular to ${\bf q}_{0}$ are gradually suppressed. Once the $\Delta_{\bf q}$ landscape fractures into four discrete maxima at $\gamma_{d}\geq 0.63\gamma_{0}$ [see the inset of Fig. 2(b)], the critical-current contour evolves into a broadly peaked unidirectional ($C_{1}$) pattern aligned with ${\bf q}_{0}$, i.e., a single preferred SC transport direction is spontaneously selected and pinned. This continuous symmetry reduction in the critical-current profile from $C_{4}$ to $C_{1}$ symmetry, marked by four, then three, and eventually a single peak, is a direct consequence of the smooth-evolution nature of the BCS-FF transition in the $d$-wave altermagnetic system.
@@
It should be emphasized that the altermagnetism-induced FF state gives rise to a significant and genuine non-reciprocal SC transport behavior, manifested by a pronounced critical-current peak around the ${\bf q}_{0}$ direction and an entirely vanishing SC response in the reverse, i.e., $J_{c}(\delta\phi=0)\neq 0$ and $J_{c}(\delta\phi=\pm\pi)=0$. This non-reciprocity is far more prominent than that in conventional FFLO states driven by isotropic Zeeman splitting [Fig. 3(b)]. In particular, deep in the FF regime, SC transport is sustained only in the vicinity of the ${\bf q}_{0}$ direction, while all other directions are suppressed, further amplifying the non-reciprocal characteristics and enhancing their relevance for practical implementation. Notably, the altermagnetism-induced FF state does not exhibit any abrupt or substantial SC gap suppression, ensuring the persistence of a robust condensate even in the highly non-reciprocal regime.
@@
Discussion.—The present study demonstrates that altermagnetic proximity offers a resilient and field-free route to realizing robust SC diode effects. While we have focused on low-temperature results to emphasize device performance, our self-consistent framework can be straightforwardly extended to finite temperatures up to $T_{c}$, ensuring relevance for realistic operating conditions. The combined features revealed here, including a non-collinear SC diode effect at weak spin splitting (i.e., in the BCS regime), pronounced SC non-reciprocity at large splitting (i.e., in the FF regime), and unidirectional SC transport deep in the FF state, along with a robust and sizable SC gap throughout the entire SC phase, make altermagnetic heterostructures promising platforms for symmetry-protected, energy-efficient SC electronics for next-generation quantum and ultra-low-dissipation circuit technologies.

### v2

@@ Abstract
Given their time-reversal-breaking yet inversion-symmetric character, and the absence of the orbital effects due to a net magnetization, altermagnets provide a natural platform for exploring their interplay with superconductivity [46, 47, 48, 49, 50, 51]. The intrinsic momentum-dependent spin polarization in altermagnets induces a mismatch between spin-up and spin-down Fermi surfaces, favoring the finite-momentum Cooper pairing even in the absence of external fields [52, 53]. Recent theoretical works [54, 55, 56, 57, 58] have investigated the stability of such pairing near the SC transition, typically by solving the linearized gap equation. A fully self-consistent microscopic analysis was performed by Hong et al. [27], who systematically compared conventional BCS pairing, chiral $p$-wave states, and FF pairing under comparable altermagnetic conditions. Their results revealed that, within a narrow but realistic range of altermagnetic strength, the FF state can become energetically favorable without requiring external magnetic fields or fine-tuned SOC.
Motivated by these developments [53, 52, 27], here we investigate directional SC transport in conventional $s$-wave SC thin films proximity-coupled [59, 60, 61, 62, 63] to a $d$-wave altermagnet, using a fully microscopic and self-consistent framework. As the altermagnetic strength increases from zero, we find that the critical current in the BCS regime evolves from being isotropic to strongly anisotropic with a fourfold ($C_{4}$) symmetry, reflecting the underlying crystalline symmetry of the altermagnetic spin texture. Thus, the directions of maximal and minimal critical currents are not antiparallel, indicating that the non-reciprocity deviates from the simple single-axis diode behavior and instead reflects a non-collinear SC diode effect.
@@
Model.—We begin with a conventional $s$-wave SC Hamiltonian, incorporating the $d$-wave altermagnetic spin splitting induced via proximity coupling [59, 60, 61] to a $d$-wave altermagnet (such as RuO2[35], ${\mathrm{Mn}}_{5}$${\mathrm{Si}}_{3}$ [31] and $\beta$-MnO2 [29, 64]). The Hamiltonian is written as [17, 65, 66, 59]
@@
Diagonalizing this Hamiltonian via Bogoliubov transformation, one obtains $H_{\bf q}=\sum_{\bf k}(E^{+}_{\bf k}\alpha_{\bf k}^{\dagger}\alpha_{\bf k}-E_{\bf k}^{-}\beta_{\bf k}^{\dagger}\beta_{\bf k}-E_{\bf k}2V_{\bf k}^{2})-\frac{\Delta_{\bf q}^{2}}{U}$. The SC free energy of the system is given by [69]
$\mathcal{F}^{s}_{\mathbf{q}}$ $=$ $\langle{H}\rangle+k_{B}T\sum_{{\bf k},\lambda=\pm,\eta=\pm}\big[f(\eta\lambda{E_{\bf k}^{\lambda}})\ln{f(\eta\lambda{E_{\bf k}^{\lambda}})}\big]$ (3)
$=$ $\sum_{{\bf k},\lambda=\pm}\big[(E_{\bf k}+\lambda\Gamma_{\bf k})f(E_{\bf k}+\lambda\Gamma_{\bf k})+k_{B}T{\ln}f(E_{\bf k}+\lambda\Gamma_{\bf k})$
$+(E_{\bf k}+\lambda\Gamma_{\bf k})\big(1-f(E_{\bf k}+\lambda\Gamma_{\bf k})\big)\big]-\sum_{\bf k}E_{\bf k}2v_{\bf k}^{2}-\frac{\Delta_{\bf q}^{2}}{U}$
$=$ $-\!\frac{1}{\beta}\sum_{\mathbf{k},\lambda=\pm}\ln\big(1\!+\!e^{-\beta\lambda E_{\mathbf{k}}^{\lambda}}\big)\!-\!\sum_{\mathbf{k}}\left(E_{\bf k}\!-\!\Xi_{\mathbf{k}}\right)\!-\!\frac{|\Delta_{\mathbf{q}}|^{2}}{U},$
where $\beta=1/(k_{B}T)$. Minimizing $\mathcal{F}^{s}_{\bf q}$ with respect to $\Delta_{\bf q}$ yields the self-consistent gap equation:
@@
where the quasiparticle energy spectrum is given by $E_{\bf k}^{\pm}=\hbar{\bf v}_{\bf k}\cdot{\bf q}+h_{\bf k}\pm{E_{\bf k}}$ with ${\bf v}_{\bf k}=\hbar{\bf k}/m$ and $E_{\bf k}=\sqrt{\Xi_{k}^{2}+\Delta_{\bf q}^{2}}$ [27, 53]. Here, $\Xi_{\bf k}=\xi_{\bf k}+2\gamma_{d}(k_{x}q_{x}-k_{y}q_{y})$, suggesting a moving-frame transformation [70], while $f(x)$ is the Fermi-Dirac distribution function, and the condensate density $2{V_{\bf k}}^{2}=1-\Xi_{\bf k}/E_{\bf k}$. The quasiparticle energies here are tilted by the finite CM momentum. As a result, at $T=0$, the integrand in the gap equation becomes nonzero only when $E_{\bf k}^{+}>0$ and $E_{\bf k}^{-}<0$. Thus, SC anomalous pairing correlations vanish in momentum-space regions where either $E_{\bf k}^{+}<0$ or $E_{\bf k}^{-}>0$, as $f(E_{\mathbf{k}}^{+})-f(E_{\mathbf{k}}^{-})\equiv 0$. These regions are referred to as unpairing regions [16, 20, 71], where electrons no longer contribute to Cooper pairing and instead behave as normal-state excitations.
@@
The equilibrium SC state is determined by self-consistently solving the gap equation and minimizing the free energy $\mathcal{F}^{s}_{\bf q}$ with respect to the pairing momentum $\bf q$}. This procedure allows us to determine the optimal CM momentum ${\bf q}_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as a function of system parameters such as the altermagnetic strength and temperature. To investigate the transport properties, we introduce an additional SC momentum ${\bf p}_{s}=\nabla_{\bf R}\theta/2-e{\bf A}$ [72, 73, 74, 75, 71, 76, 65, 17, 77, 78], corresponding to a uniform phase gradient, superflow, or external vector potential. The resulting supercurrent is given by ${\bf j}_{s}=(e/m)n_{s}{\bf p}_{s}$ [72, 73, 74, 75, 71, 76, 65, 79], where $n_{s}$ is the superfluid density [80], which approaches the total electron density $n$ in the low-temperature limit [76, 68, 20, 72, 17, 81, 82]. To determine the critical current, we gradually increase the magnitude of ${\bf p}_{s}$ along a fixed direction and solve the gap equation self-consistently under the total momentum ${\bf q}_{0}+{\bf p}_{s}$. The SC gap collapses at a critical value ${\bf p}_{s,c}$, signaling the pair-breaking threshold and marking the onset of dissipation.
@@
For comparison, we also consider a conventional FFLO model with an isotropic spin splitting arising from ferromagnetic proximity, modeled as $h_{\bf k}=\gamma_{s}k^{2}$. This serves to isolate the role of the anisotropic, crystal-symmetry-protected spin splitting characteristic of altermagnets.
@@
Figure 2: Simulated order-parameter (OP) results for the optimal CM momentum $q_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as functions of the spin-splitting strength for (a) ferromagnetic and (b) $d$-wave altermagnetic proximity. Insets: the $\Delta_{\bf q}$ landscape plotted in $q_{x}$$q_{y}$ plane, which reflects the inverse of the free-energy profile. The plotted $q$ range extends up to $0.8q_{0}$ in each direction. T=0.1T=0.1~K. Normalized parameters used here are: the zero-splitting gap $\Delta_{0}=2$~meV, coherence length $\xi_{0}=v_{F}/\Delta_{0}$, and splitting unit $\gamma_{0}=\Delta_{0}/k_{F}^{2}$.
@@
Phase diagram.— As a specific example, we consider a typical SC thin film, NbN, which has a SC gap $\Delta_{0}\sim 2~\text{meV}$ at $T\sim 0$ and a critical SC temperature $T_{c}\sim 12$-15 K [68, 83, 84]. Due to the Cooper instability of interacting fermions underlying the BCS theory [17, 66, 65, 85], the momentum summation for the pairing electrons is restricted such that the spin-up electron with momentum $\bf k+q$} and the spin-down electron with momentum $\bf-k+q$} both lie within their respective Debye momentum shells around the Fermi surface, enforced by the spin-dependent condition $|\xi_{s{\bf k+q}}+sh_{s{\bf k+q}}|\leq\omega_{D}$. The CM momentum dependence of $\Delta_{\bf q}$ directly reflects the inverse of the free-energy landscape since $\mathcal{F}^{s}_{\mathbf{q}}-\mathcal{F}^{n}_{\mathbf{q}}\propto-\Delta_{\bf q}^{2}$, offering an efficient means for following analysis.
@@
In the BCS phase at the ferromagnetic-proximity case [Fig. 3(a)], $J_{c}(\delta\phi)$ is fully isotropic and decreases gradually with increasing $\gamma_{s}$, as expected. Upon entering the FF phase via a first-order transition [Fig. 3(b)], the critical-current contour becomes strongly anisotropic, and a clear forward–backward asymmetry emerges on the ${\bf q}_{0}$ axis, i.e., $J_{c}(\delta\phi=0){\neq}J_{c}(\delta\phi=\pm\pi)$, manifesting as a SC-diode effect in agreement with earlier prediction [7]. Interestingly, the largest critical currents do not occur precisely at $\delta\phi=\pi$, but instead appear in the range $\delta\phi\approx 0.75$$0.5\pi$, while the minimum lies at $\delta\phi=0$. This shift, which has been overlooked in previous studies, has a geometric origin arising from the circular symmetry of the $\Delta_{\bf q}$ shell in the FF state. Specifically, in the conventional FF state, the gap landscape $\Delta_{\bf q}$ [e.g., the region with a nonzero $\Delta_{\bf q}$ in the inset of Fig. 2(a) for $\gamma_{s}=0.45\gamma_{0}$] is nearly degenerate along a circular shell in $\bf q$}-space, with the optimal pairing momentum ${\bf q}_{0}$ located near its outer edge. As a result, a supercurrent applied parallel to ${\bf q}_{0}$ quickly drives the total momentum to the edge of this shell and triggers the gap to collapse, yielding the minimum $J_{c}$. A current slightly off the antiparallel direction probes more stable segments of the shell, thereby resulting in the maximum $J_{c}$.
@@
For an altermagnetic exchange field, the evolution of the critical current exhibits qualitatively distinct behavior. As the splitting increases from zero, the BCS state [Fig. 3(c)] gradually loses its isotropy, and the angular profile of $J_{c}(\delta\phi)$ starts to develop the pronounced oscillations. With increasing $\gamma_{d}$, the critical current along the spin-splitting axes (i.e., the $x$ and $y$ directions: $\delta\phi=\pm\pi/4$ and $\pm 3\pi/4$) is progressively suppressed, while the critical current along the nodal directions of the splitting (i.e., the diagonal directions: $\delta\phi=0,\pi$ and $\pm\pi/2$) is enhanced. As a result, $J_{c}(\delta\phi)$ acquires a pronounced fourfold ($C_{4}$) anisotropy, directly reflecting the $d$-wave character of the altermagnetic spin texture and corresponding to the $C_{4}$-symmetric, square-like shape of the $\Delta_{\bf q}$ landscape below the critical value [see the insets of Fig. 2(b)]. Thus, the maxima of $J_{c}(\delta\phi)$ occur along the diagonal directions, while the minima lie along $x$ and $y$ axes. This nontrivial angular misalignment between the directions of maximal and minimal critical currents gives rise to a non-collinear SC diode effect, where the non-reciprocity of the supercurrent is not aligned with a single transport direction, but instead emerges from distinct crystal axes. This non-collinear SC diode effect is robust, as it originates from the crystal-symmetry-protected nature of the altermagnetic spin splitting, enabling intrinsically stable and symmetry-guided rectification of supercurrents.
@@
With further increase of $\gamma_{d}$, following the onset of the FF phase, a finite CM momentum ${\bf q}_{0}$ develops along one of the nodal directions of the altermagnetic spin splitting, and directly breaks SC transport along the opposite direction. This asymmetry occurs because ${\bf q}_{0}$ resides near the inner boundary of the gap landscape $\Delta_{\bf q}$, making the system particularly sensitive to supercurrents applied in the ${\bf q}_{0}$ direction, which easily push the total momentum beyond the de-pairing threshold and trigger SC breakdown. As a result, only three residual peaks remain in the critical-current profile, aligned with the other nodal directions ($\delta\phi=0,\pm\pi/2$), since in this early FF regime, the $\Delta_{\bf q}$ landscape still spans a continuous connected region in $\bf q$} space [see the inset of Fig. 2(b)], although its global maximum has already shifted away from the origin to ${\bf q}_{0}$. Notably, as $\gamma_{d}$ increases further, the critical-current peaks perpendicular to ${\bf q}_{0}$ are gradually suppressed. Once the $\Delta_{\bf q}$ landscape fractures into four discrete maxima at $\gamma_{d}\geq 0.63\gamma_{0}$ [see the inset of Fig. 2(b)], the critical-current contour evolves into a broadly peaked unidirectional ($C_{1}$) pattern aligned with ${\bf q}_{0}$, i.e., a single preferred SC transport direction is spontaneously selected and pinned. This continuous symmetry reduction in the critical-current profile from $C_{4}$ to $C_{1}$ symmetry, marked by four, then three, and eventually a single peak, is a direct consequence of the smooth-evolution nature of the BCS-FF transition in the $d$-wave altermagnetic system.
@@
It should be emphasized that the altermagnetism-induced FF state gives rise to a significant genuine and nearly ideal non-reciprocal SC transport behavior, manifested by a pronounced critical-current peak around the ${\bf q}_{0}$ direction and an entirely vanishing SC response in the reverse, i.e., $J_{c}(\delta\phi=0)\neq 0$ and $J_{c}(\delta\phi=\pm\pi)=0$. This non-reciprocity is far more prominent than that in conventional FFLO states driven by isotropic splitting [Fig. 3(b)]. In particular, deep in the FF regime, SC transport is sustained only in the vicinity of the ${\bf q}_{0}$ direction, while all other directions are suppressed, further amplifying the non-reciprocal characteristics and enhancing their relevance for practical implementation. Notably, the altermagnetism-induced FF state does not exhibit any abrupt or substantial SC gap suppression, ensuring the persistence of a robust condensate even in the highly non-reciprocal regime.
Figure 4: Temperature–splitting phase diagram of the (a) SC gap and (b) the optimal CM momentum. (c) Overall phase diagram, where the BCS, normal, and FF phases are indicated in blue, red, and green, respectively.
@@
Discussion.—The present study demonstrates that altermagnetic proximity offers a resilient and field-free route to realizing robust SC diode effects. The combined features revealed, including a non-collinear SC diode effect at weak spin splitting (i.e., in the BCS regime), pronounced SC non-reciprocity at large splitting (i.e., in the FF regime), and unidirectional SC transport deep in the FF state, along with a robust and sizable SC gap throughout the entire SC phase, make altermagnetic heterostructures promising platforms for symmetry-protected, energy-efficient SC electronics for next-generation quantum and ultra-low-dissipation circuit technologies. Physically, the gap equation at $q=0$ is an even function of $h_{\bf k}$ and remains invariant under $C_{4}$. One therefore expects that the SC instability and the resulting critical currents are $C_{4}$-symmetric. After the system develops a finite pairing momentum $\mathbf{q}$, the spontaneously selected direction of $\mathbf{q}$ reduces the rotational symmetry from $C_{4}$ to $C_{1}$, giving rise to the diode behavior.
While we have focused on low-temperature results to emphasize device performance, the framework can be straightforwardly extended to finite temperatures up to the critical temperature $T_{c}$. A complete temperature–splitting phase diagram of the order parameters (the SC gap and the optimal CM momentum) is presented in Fig. 4. As shown in the figure, with increasing temperature, (I) for $\gamma_{d}\in(0,0.2\gamma_{0})$, the system undergoes a second-order transition from the BCS to the normal state, characterized by a continuous suppression of the SC gap; (II) for $\gamma_{d}\in(0.2\gamma_{0},0.5\gamma_{0})$, the transition becomes first-order, marked by a discontinuous collapse of the gap, indicative of phase competition or metastability; (III) for $\gamma_{d}\in(0.5\gamma_{0},0.66\gamma_{0})$, the system first undergoes a second-order transition from the BCS to the finite-momentum FF state, followed by a first-order transition from the FF to the normal state, suggesting enhanced sensitivity to thermal fluctuations in the finite-momentum condensate. A tri-critical Lifshitz point appears at approximately $T\approx 7$ K and $\gamma_{d}\approx 0.5\gamma_{0}$, where three distinct phases (BCS, FF, and normal) coexist. This point simultaneously marks (i) the onset of a finite pairing momentum, (ii) the change in the nature of the SC transition, and (iii) a topology change in the phase diagram. Such a Lifshitz point reflects a qualitative restructuring of the SC ground state, and may manifest experimentally as a nonanalyticity in thermodynamic quantities or anomalous transport behavior near the transition[87, 88]. Moreover, in Supplement Materials [89] (see also references [90, 91] therein), we have also derived the emergence of spin-triplet pairing correlations [59, 92, 20] which may lead to novel phenomena such as an anomalous paramagnetic Meissner effect, as introduced in Refs. [93, 63] and experimentally observed in Ref. [94], and cause the magnetoelectric effect [20, 95].